OpenStudy (anonymous):

Mario Agnello is opening a new pizza shop in town. He plans to offer a 12 -inch diameter pizza and an 18-inch diameter pizza. He has set the price of a 12-inch cheese pizza at $8.00 , based on the amount of ingredients needed to cover that size crust. If Mario would like to keep the pricing proportional, what should he charge for an -inch cheese pizza? Show and explain your solution.

OpenStudy (ghazi):

$8 dollars for 12 inch $1 for 12/8 inch= 1.5 inch or else you can say... 12 inch for $8 1 inch for $ 8/12 = $ 0.667

OpenStudy (ghazi):

@amanduh is this clear?

OpenStudy (anonymous):

start by finding the area of a 12 in-diameter and a 18 in-diameter pizza \[A=\frac{\pi}{4}d^2=\pi r^2\] then figure out the proportion

OpenStudy (anonymous):

\[A_{12 inch}= 36\pi\] \[A_{18 inch}= 81\pi\] \[\frac{price of 12 inch}{area of 12 inch}=\frac{price of 18 inch}{area of 18 inch}\] the price of a 12 in diameter pizza is 8 dollars

OpenStudy (anonymous):

Mkay thank you guys... sorry for the really late response my internet was down